Four particle are to be distributed in 5 states. Calculate the number of ways in which this distribution can be done if the particles obey:


1) M-B statistics 2)B-E statistics.

1
Expert's answer
2022-03-16T18:29:52-0400

Solution;

M=4

N=5

"W=(\displaystyle_5^4)=\frac{4!}{5!(5-4)!}"

"W=\frac{1}{5}"